1. Solve the initial value problem =[71]e, 20 = [2] Describe the behavior of the solution as t + 0.
Answer #1

x=( ) 5 tacoma (-1). The charactens He eqs IA - XI] 30 15-2 Il-nio leo judi b ogati 13 l-xl (5-2 (1-1) +3=0 5-50-X+W²+350 xx, = ($). for a=4 (A-4I) X2 = 0 Ġ 3)(!) =(:) 3 R2 R2-3R, Cool 2 2 ) = (8) . 21-nico x,=42 since the rank of this matna is als